# Space-Efficient Hierholzer: Eulerian Cycles in $\mathrm{O}(m)$ Time and $\mathrm{O}(n)$ Space

### v1

@@ 1 Introduction
An Eulerian cycle in a graph is a closed walk that traverses every edge exactly once. Euler’s study of their existence and computation is often cited as the birth of modern graph theory. Apart from historical significance, efficiently computing Eulerian cycles has applications, e.g., in genome assembly [6], routing problems [5], and 3D printing [8].
@@
In this paper, we present a simple variant of Hierholzer’s algorithm that uses only $\mathrm{O}(n\lg m)$ bits of working memory and runs in optimal $\mathrm{O}(n+m)$ time. For dense graphs or multigraphs with many parallel edges, this is a substantial saving. Note that this space can neither hold the entire input nor the entire output of the problem; we treat the input graph as given in read-only memory and produce the Eulerian cycle in order to a write-only stream. In particular, our algorithm would be suitable for an application that directly consumes and uses edges of the Eulerian cycle as they are computed, potentially without ever storing the entire cycle. By working memory, we mean the extra space occupied by auxiliary data structures during the computation.
To our knowledge, we present the first algorithm with working space $\mathrm{O}(m\lg n)$ bits that runs in linear time and explicitly produces an Eulerian cycle as its output. Our algorithm is also usable in rule-based models of computation, such as graph-transformation languages like GP 2 [7], which do not natively support stacks or recursion. This is not known to be true for other standard efficient implementations of Eulerian cycle algorithms.
@@ 1.2 Typical Implementations
More amenable to producing the output directly in the correct order is the approach to treat Hierholzer’s algorithm as an edge-centric depth-first-search. We follow the basic steps of a graph traversal, but mark edges rather than vertices. So we only backtrack from a vertex when all of its incident edges have been traversed. When this happens for the first time, the last edge traversed to this dead end was the edge closing the first cycle of Hierholzer’s algorithm, and can thus be output now as the last edge of the Eulerian cycle. Iteratively, the same is true for the next dead end vertex.222 See, e.g., slide 33 of these lecture notes: https://www.wild-inter.net/teaching/ea/09-graph-algorithms.pdf#page=40. Both the stack and the marking of edges require $\Theta(m\lg m)$ bits of working space in the worst case. A minor twist of this approach, where edges are not marked, but removed from a copy of the input graph, is used, e.g., in NetworkX [1], a popular graph framework for Python. The copy, of course, also uses $\Omega(m\lg n)$ bits of space.
@@ 1.5 Related Work
Glazik et al. [3] present a streaming algorithm for Eulerian cycles. While they also achieve $\mathrm{O}(n\lg n)$-space and in a much more restrictive model, it is not clear if their algorithm can be implemented to run in linear total time, and they construct the cycle implicitly via a successor function. Their algorithm is arguably more complicated than Hierholzer’s algorithm.
@@ 2 Preliminaries
We write $[n]$ for $\{1,2,\dots,n\}$. Throughout the paper, $G=(V,E)$ denotes a finite directed multigraph, where $V$ is the set of vertices and $E$ is the multiset of directed edges, with $|V|=n$ and $|E|=m$. Loop edges are allowed. We assume w.l.o.g. that $V=[n]$. We denote by $d^{+}(v)$ and $d^{-}(v)$ the out-degree and in-degree of vertex $v$, respectively. The $i$th out-neighbour of a vertex $v$ is denoted by $\Gamma^{+}(v,i)$, and its $i$th in-neighbour by $\Gamma^{-}(v,i)$, assuming arbitrary but fixed orderings of incoming and outgoing edges for each vertex. We concisely write $uv$ for a (directed) edge from $u$ to $v$.
@@
We assume that basic arithmetic and memory operations on $\lg m$-bit integers take constant time, and that the in- and out-degree functions as well as neighbour queries can be evaluated in constant time per operation.
@@ 3.1 Conceptual Algorithm
Let $G$ be a directed Eulerian graph, and let $v_{0}$ be the starting vertex (it may be chosen arbitrarily). Unlike the classical approach, our algorithm traverses the graph by following incoming edges rather than outgoing ones; that is, we perform a reversed traversal of the graph. This reversal allows the algorithm to compute the tour incrementally, in the correct order, without needing to store the entire tour and reverse it at the end.
@@ 3.2 Space-Efficient Implementation
[…]
18:Write(uv)(uv)
[…]
@@
When a vertex $v$ is visited for the first time (via an outgoing edge $vu$), the algorithm records $u$ as $v$’s predecessor by setting $B[v]\leftarrow u$; this is equivalent to marking $uv$ Red. Backtracking begins when a vertex $u$ has no unexplored incoming (Black) edges, which occurs when $\texttt{next}[u]>d^{-}(u)$. At that point, the algorithm starts following outgoing edges from $u$. To ensure the Red edge is used last, it checks whether the next candidate edge (indexed by $\texttt{next}[u]-d^{-}(u)$) leads to $B[u]$. If so, and if it has not already been skipped, the algorithm skips this edge. This guarantees that all non-Red (i.e. Green) edges are backtracked over first. [^4] Since the graph may contain parallel edges, the target vertex $v=B[u]$ could appear multiple times in the multiset $\{\Gamma^{+}(u,i):i\in[d^{+}(u)]\}$. Hence we use skipped-array to only skip the first edge $uv$. Finally, after all other outgoing edges have been used (i.e. when $\texttt{next}[u]>d^{+}(u)+d^{-}(u)=d(u)$), the Red edge becomes the only option and is followed. As the algorithm backtracks from vertices, each traversed edge is immediately written to the output.
@@ 4.1 Proof of Correctness
  2. $2.$
    $x\neq a.$ Before entering $x$ from $b$, we have $d^{+}(x,\textsc{Black})=d^{-}(x,\textsc{Black})$; upon entering it, we have $d^{+}(x,\textsc{Black})+1=d^{-}(x,\textsc{Black})$, $d^{+}(b,\textsc{Black})=d^{-}(b,\textsc{Black})$, $d^{+}(a,\textsc{Black})=d^{-}(a,\textsc{Black})+1$, and $x$ is the new current vertex.
@@
Statement 1 of Lemma 1 holds. Before backtracking, we have $d^{+}(v,\textsc{Black})=d^{-}(v,\textsc{Black})$ for all $v\in V$. Since there is exactly one needle vertex, then it must be the current vertex $u$ by Definition 2, which is in contradiction with the assumption that the current vertex is not the needle vertex.
@@
A consequence of Lemma 2 is that backtracking occurs for the first time at $v_{0}$, and then makes a vertex incident on $v_{0}$ the new needle vertex. Observe that this process indeed forms a trail of Dashed edges in the graph. We now show that this trail is closed at termination, i.e. forms a cycle where $v_{0}$ is included. To do so, it suffices to show that the algorithm terminates and that $v_{0}$ is the last current vertex. To prove termination, we fix the cost function $c:M\to\mathbb{N}\cup\{0\}$ s.t. $M=\{\textsc{Black},~\textsc{Red},~\textsc{Green},~\textsc{Dashed}\}$, $c(\textsc{Black})=3$, $c(\textsc{Red})=2$, $c(\textsc{Green})=2$, and $c(\textsc{Dashed})=0$.
@@
Observe that, by the definition of an Eulerian graph, entering a non-starting vertex via an outgoing Black edge (which is then marked Red or Green) guarantees that there is a corresponding incoming Black edge that will eventually be used to exit the vertex and marked accordingly. Similarly, entering a non-starting vertex via a Red or Green outgoing edge (which is then marked Dashed) ensures that there is a corresponding incoming Red or Green edge that will eventually be used to exit the vertex and also be marked Dashed. Therefore, at every step of the algorithm, if $u$ is the current vertex, the following conditions hold:
@@
A Black edge is incident on a vertex $v\in T$. If $v$ has a Black incoming edge, this implies that the algorithm backtracked from or terminated at $v$ despite there being a Black incoming edge to traverse, which is a contradiction. If, however, the edge is outgoing from $v$, this implies that there is a Black incoming edge at $v$ by the conditions of the input, and the previous subcase applies.
@@
First, observe that the Red edges marked by Space-Efficient-Hierholzer form a tree rooted at $v_{0}$, the starting vertex, and hence $d^{+}(v,\textsc{Red})\leq 1$ for any vertex $v$. Pick an arbitrary set $S_{i}$: for each $w\in S_{i}$, we have $d^{+}(w,\textsc{Red})+d^{+}(w,\textsc{Green})=d^{-}(w,\textsc{Red})+d^{-}(w,\textsc{Green})>0$. If $v_{0}\in S_{i}$, then we are done since $d^{+}(v_{0},\textsc{Red})=0$ and thus $d^{+}(v_{0},\textsc{Green})\geq 1$. Now, towards a contradiction, suppose that we have $d^{+}(w,\textsc{Red})=1$ for all $w\in S_{i}$ (recall that this measure cannot be greater than $1$), and $v_{0}\not\in S_{i}$. Since $H_{i}$ does not share vertices with other components in $\biguplus_{j\in[k]\setminus\{i\}}H_{j}$, there is no path from a vertex in $H_{i}\setminus S_{i}$ to $v_{0}$ that does not include an outgoing edge incident on a vertex in $S_{i}$. Thus, the assumption entails that the starting vertex (i.e. the root of the tree of Red edges) $v_{0}\in H_{i}\setminus S_{i}$, which contradicts our assumption that $v_{0}\in T$ since $T\cap(H_{i}\setminus S_{i})=\emptyset$.
@@
The statement seamlessly follows from Corollary 2. ∎
@@ 4.2 Space and Time Complexity
The overall complexity thus is $\mathrm{O}(n+m)\in\mathrm{O}(m)$ since, in any Eulerian graph of $2$ or more vertices, $n\leq m$. ∎
@@ 5 Conclusion
Our approach focuses on directed graphs, but the underlying principle can also be applied to undirected graphs. A nuisance not present in directed graphs is that an edge might have been traversed in the opposite direction when we now consider it, so we have to detect this. When the undirected graph is simple (i.e. no parallel edges) and that the adjacency lists are sorted – that is, for all $i<j\in[d(v)]$, we have $\Gamma(v,i)<\Gamma(v,j)$ for all $v\in V$ – we can determine whether an edge $uv$ incident at $u$ has already been traversed via $\Gamma(v,\texttt{next}[v])\geq u$.

### v2

@@ 1 Introduction
An Eulerian cycle in a graph is a closed walk that traverses every edge exactly once. A graph is called Eulerian if it contains an Eulerian cycle. Euler’s study of the existence of Eulerian cycles and their computation is often cited as the birth of modern graph theory. Apart from historical significance, efficiently computing them has applications in, e.g., genome assembly [7], routing problems [6], and 3D printing [9].
@@
In this paper, we present a simple variant of Hierholzer’s algorithm that uses only $\mathrm{O}(n\lg m)$ bits of working memory and runs in optimal $\mathrm{O}(n+m)$ time. For dense graphs or multigraphs with many parallel edges, this is a substantial saving. Note that this space can neither hold the entire input nor the entire output of the problem; we treat the input graph as given in read-only memory and produce the Eulerian cycle in order to a write-only stream. In particular, our algorithm would be suitable for an application that directly consumes and uses edges of the Eulerian cycle as they are computed, potentially without ever storing the entire cycle. By working memory, we mean the extra space occupied by auxiliary data structures during the computation. Our algorithm relies on an assumption about the graph representation, which is satisfied in typical adjacency-list implementations: for directed graphs, we require that we can iterate over incoming and outgoing edges; for undirected graphs, we (instead) require a consistent ordering of vertices across all adjacency lists.
To our knowledge, we present the first algorithm with working space $\mathrm{O}(n\lg m)$ bits that runs in linear time and explicitly produces an Eulerian cycle as its output. Our algorithm is also usable in rule-based models of computation, such as graph-transformation languages like GP 2 [8], which do not natively support stacks or recursion. This is not known to be true for other standard efficient implementations of Eulerian cycle algorithms.
@@ 1.2 Typical Implementations
More amenable to producing the output directly in the correct order is the approach to treat Hierholzer’s algorithm as an edge-centric depth-first-search. We follow the basic steps of a graph traversal, but mark edges rather than vertices. So we only backtrack from a vertex when all of its incident edges have been traversed. When this happens for the first time, the last edge traversed to this dead end was the edge closing the first cycle of Hierholzer’s algorithm, and can thus be output now as the last edge of the Eulerian cycle. Iteratively, the same is true for the next dead end vertex. [^2] See, e.g., slide 33 of these lecture notes: https://www.wild-inter.net/teaching/ea/09-graph-algorithms.pdf#page=40. Both the stack and the marking of edges require $\Theta(m\lg m)$ bits of working space in the worst case. A minor twist of this approach, where edges are not marked, but removed from a copy of the input graph, is used, e.g., in NetworkX [1], a popular graph framework for Python. The copy, of course, also uses $\Omega(m\lg n)$ bits of space.
@@ 1.5 Related Work
Glazik, Schiemann, and Srivastav [3] present a streaming algorithm for Eulerian cycles. While they also achieve $\mathrm{O}(n\lg n)$-space and in a much more restrictive model, it is not clear if their algorithm can be implemented to run in linear total time, and they construct the cycle implicitly via a successor function. Their algorithm is arguably more complicated than Hierholzer’s algorithm.
Hagerup, Kammer, and Laudahn [4] proposed an algorithm using $\mathrm{O}(n+m)$ time and $\mathrm{O}(n+m)$ bits of space for computing Eulerian cycles. They do not need strong assumptions regarding the graph data representation, but focus on undirected graphs only. They build a nontrivial dynamic data structure in each vertex that stores which pairs of edges are adjacent in the Eulerian cycle. Their algorithm is considerably more involved than the typical implementations of Hierholzer’s algorithm using $\mathrm{O}(n+m)$ words of space.
@@ 2 Preliminaries
We write $[n]$ for $\{1,2,\dots,n\}$. Throughout the paper, $G=(V,E)$ denotes a finite directed multigraph, where $V$ is the set of vertices and $E$ is the multiset of directed edges, with $|V|=n$ and $|E|=m$. Loop edges are allowed. We assume without loss of generality that $V=[n]$. We denote by $d^{+}(v)$ and $d^{-}(v)$ the out-degree and in-degree of vertex $v$, respectively. The $i$th out-neighbour of a vertex $v$ is denoted by $\Gamma^{+}(v,i)$, and its $i$th in-neighbour by $\Gamma^{-}(v,i)$, assuming arbitrary but fixed orderings of incoming and outgoing edges for each vertex. We concisely write $uv$ for a (directed) edge from $u$ to $v$.
@@
A trail in a graph is a walk in which all edges are distinct. A closed trail is a trail whose first and last vertices coincide.
We assume that basic arithmetic and memory operations on $\lg m$-bit integers take constant time, and that the in- and out-degree functions as well as in- and out-neighbour queries can be evaluated in constant time per operation.
@@ 3.1 Conceptual Algorithm
Let $G$ be a directed Eulerian graph, and let $v_{0}$ be the starting vertex (it may be chosen arbitrarily). Unlike the classical approach, our algorithm traverses the graph by following incoming edges rather than outgoing ones; that is, we perform a reverse traversal of the graph. This reversal allows the algorithm to compute the tour incrementally, in the correct order, without needing to store the entire tour and reverse it at the end.
@@ 3.2 Space-Efficient Implementation
1: $\texttt{next}[1..n],~\texttt{visited}[1..n],~\texttt{skipped}[1..n],~B[1..n]\leftarrow[0]^{n}$
2: $c\leftarrow 0$
3: $v_{0}\leftarrow k\in[n]$ $\triangleright$ Pick an arbitrary starting vertex.
4: $u\leftarrow v_{0}$ $\triangleright$ Initially, the current vertex is the starting one.
5: $\texttt{visited}[u]\leftarrow 1$
6: while $c<m$ do $\triangleright$ We loop as long as there are still edges to write.
7:   $\texttt{next}[u]\leftarrow\texttt{next}[u]+1$
8:   $i\leftarrow\texttt{next}[u]$
9:   if $i\leq d^{-}(u)$ then $\triangleright$ Reverse traversal phase.
10:    $v\leftarrow\Gamma^{-}(u,i)$
11:    if $\texttt{visited}[v]=0$ then
12:      if $v\neq v_{0}$ then
13:       $B[v]\leftarrow u$
14:      $\texttt{visited}[v]\leftarrow 1$
15:    $u\leftarrow v$
16:   else $\triangleright$ Backtracking phase.
17:    $i\leftarrow\texttt{next}[u]-d^{-}(u)$
18:    if $\Gamma^{+}(u,i)=B[u]$ and $\texttt{skipped}[u]=0$ then
19:      $\texttt{skipped}[u]\leftarrow 1$
20:      $\texttt{next}[u]\leftarrow\texttt{next}[u]+1$
21:      $i\leftarrow i+1$
22:    if $i>d^{+}(u)$ then
23:      $v\leftarrow B[u]$
24:    else
25:      $v\leftarrow\Gamma^{+}(u,i)$
26:    Write(uv)(uv)
27:    $c\leftarrow c+1$
28:    $u\leftarrow v$
@@
When a vertex $v$ is visited for the first time (via an outgoing edge $vu$), the algorithm records $u$ as $v$’s predecessor by setting $B[v]\leftarrow u$; this is equivalent to marking $uv$ Red. Backtracking begins when a vertex $u$ has no unexplored incoming (Black) edges, which occurs when $\texttt{next}[u]>d^{-}(u)$ (i.e. when $\texttt{next}[u]$ is out of the bounds of $d^{-}(u)$). At that point, the algorithm starts following outgoing edges from $u$. To ensure the Red edge is used last, it checks whether the next candidate edge (indexed by $\texttt{next}[u]-d^{-}(u)$) leads to $B[u]$. If so, and if it has not already been skipped, the algorithm skips this edge. This guarantees that all non-Red (i.e. Green) edges are backtracked over first. [^4] Since the graph may contain parallel edges, the target vertex $v=B[u]$ could appear multiple times in the multiset $\{\Gamma^{+}(u,i):i\in[d^{+}(u)]\}$. Hence we use skipped-array to only skip the first edge $uv$. Finally, after all other outgoing edges have been used (i.e. when $\texttt{next}[u]>d^{+}(u)+d^{-}(u)=d(u)$), the Red edge becomes the only option and is followed. As the algorithm backtracks from vertices, each traversed edge is immediately written to the output.
@@ 4.1 Proof of Correctness
Subcase 2.2.
$x\neq a.$ Before entering $x$ from $b$, we have $d^{+}(x,\textsc{Black})=d^{-}(x,\textsc{Black})$; upon entering it, we now have $d^{+}(x,\textsc{Black})+1=d^{-}(x,\textsc{Black})$, $d^{+}(b,\textsc{Black})=d^{-}(b,\textsc{Black})$, $d^{+}(a,\textsc{Black})=d^{-}(a,\textsc{Black})+1$, and $x$ is the new current vertex. ∎
@@
Statement 1 of Lemma 1 holds. For all $v\in V$, we have $d^{+}(v,\textsc{Black})=d^{-}(v,\textsc{Black})$ before backtracking. Since there is exactly one needle vertex, then it must be the current vertex $u$ by Definition 2, which is in contradiction with the assumption that the current vertex is not the needle vertex.
@@
This completes the proof. ∎
Lemma 3.
The needle vertex changes only upon backtracking.
Proof.
Towards a contradiction, suppose otherwise; that is, assume that a forward step into $v$ makes $v$ the needle vertex. Before this forward step, we had $d^{+}(v,\textsc{Black})=d^{-}(v,\textsc{Black})$. After the step, however, $d^{+}(v,\textsc{Black})+1=d^{-}(v,\textsc{Black})$. Since $v$ is now the current vertex and $d^{+}(v,\textsc{Black})\neq d^{-}(v,\textsc{Black})$, this contradicts Definition 2. ∎
Remark 1.
A consequence of Lemmata 2 and 3 is that backtracking occurs for the first time at $v_{0}$ (the first needle vertex), and then makes a vertex incident on $v_{0}$ the new needle vertex. Observe that this process indeed forms a trail of Dashed edges in the graph.
We now show that this trail is closed at termination, i.e. forms a cycle where $v_{0}$ is included. To do so, it suffices to show that the algorithm terminates and that $v_{0}$ is the last current vertex. To prove termination, we fix the cost function $c:M\to\mathbb{N}\cup\{0\}$ s.t. $M=\{\textsc{Black},~\textsc{Red},~\textsc{Green},~\textsc{Dashed}\}$, $c(\textsc{Black})=3$, $c(\textsc{Red})=2$, $c(\textsc{Green})=2$, and $c(\textsc{Dashed})=0$.
@@
By Remark 1, the algorithm first backtracks from $v_{0}$. Observe that, by the definition of an Eulerian graph, entering a non-starting vertex via an outgoing Black edge (which is then marked Red or Green) guarantees that there is a corresponding incoming Black edge that will eventually be used to exit the vertex and marked accordingly. Similarly, entering a non-starting vertex via a Red or Green outgoing edge (which is then marked Dashed) ensures that there is a corresponding incoming Red or Green edge that will eventually be used to exit the vertex and also be marked Dashed. Therefore, at every step of the algorithm, if $u$ is the current vertex, the following conditions hold:
@@
A Black edge is incident on a vertex $v\in T$. If $v$ has a Black incoming edge, this implies that the algorithm backtracked from or terminated at $v$ despite there being a Black incoming edge to traverse, which is a contradiction. If, on the other hand, the Black edge is outgoing from $v$, then by Lemma 1, there must also be an incoming Black edge at $v$. Indeed, Statement 1 of the lemma applies, since the current vertex cannot possess incoming Black edges without violating the algorithm’s definition (i.e. backtracking would have occurred instead of termination). Therefore, the previous subcase applies.
@@
First, observe that the Red edges marked by Space-Efficient-Hierholzer form a tree rooted at $v_{0}$, the starting vertex; hence $d^{+}(v,\textsc{Red})\leq 1$ for any vertex $v$. Pick an arbitrary set $S_{i}$: using the invariant from the proof of Lemma 5, for each $w\in S_{i}$, we have $d^{+}(w,\textsc{Red})+d^{+}(w,\textsc{Green})=d^{-}(w,\textsc{Red})+d^{-}(w,\textsc{Green})>0$. If $v_{0}\in S_{i}$, then we are done since $d^{+}(v_{0},\textsc{Red})=0$ and thus $d^{+}(v_{0},\textsc{Green})\geq 1$. Now, towards a contradiction, suppose that we have $d^{+}(w,\textsc{Red})=1$ for all $w\in S_{i}$ (recall that this measure cannot be greater than $1$), and $v_{0}\not\in S_{i}$. Since $H_{i}$ does not share vertices with other components in $\biguplus_{j\in[k]\setminus\{i\}}H_{j}$, there is no path from a vertex in $H_{i}\setminus S_{i}$ to $v_{0}$ that does not include an outgoing edge incident on a vertex in $S_{i}$. Thus, the assumption entails that the starting vertex (i.e. the root of the tree of Red edges) $v_{0}\in H_{i}\setminus S_{i}$, which contradicts our assumption that $v_{0}\in T$ since $T\cap(H_{i}\setminus S_{i})=\emptyset$.
@@
By Lemma 4, the algorithm terminates after a finite number of steps. Lemma 5 establishes that, upon termination, the set of Dashed edges forms a closed trail containing the starting vertex $v_{0}$. Moreover, Lemma 3 and Lemma 2 together ensure that a new Dashed edge is created exactly when the traversal backtracks from the current needle vertex, and that the next needle vertex is incident to the previously Dashed one, thus producing the trail in a correct order. Finally, Corollary 2 implies that all edges are marked Dashed upon termination, which ensures that the closed trail includes every edge exactly once. ∎
@@ 4.2 Space and Time Complexity
The overall complexity is $\mathrm{O}(n+m)\in\mathrm{O}(m)$ since, in any Eulerian graph of $2$ or more vertices, $n\leq m$. ∎
@@ 5 Conclusion
Our approach focuses on directed graphs, but the underlying principle can also be applied to undirected graphs. A nuisance not present in directed graphs is that an edge might have been traversed in the opposite direction when we now consider it, so we have to detect this. When the undirected graph is simple (i.e. no parallel edges) and the adjacency lists are sorted (that is, for all $i<j\in[d(v)]$, we have $\Gamma(v,i)<\Gamma(v,j)$ for all $v\in V$), we can determine whether an edge $uv$ incident at $u$ has already been traversed via $\Gamma(v,\texttt{next}[v])\geq u$.